Status Report

Hubble Space Telescope Daily Report #2722 10/05/00

By SpaceRef Editor
October 5, 2000
Filed under

HUBBLE SPACE TELESCOPE
 
  DAILY REPORT #2722
 
PERIOD COVERED:  0000Z (UTC) 10/04/00 – 0000Z (UTC) 10/05/00
 
Daily Status Report as of 279/0000Z
 
1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED:
 
    1.1 Completed STIS/CCD/MA2 7316 (Saturn’s Satellites: Alteration of the Surface by Ion Irradiation on Enceladus and Tethys)
 
        The Space Telescope Imaging Spectrograph (CCD and MA2) was used to make observations of Saturn’s moon Tethys.  Saturn’s medium-sized icy satellites display a diverse range of surface alteration caused by one or more external processes.  A wavelength calibration image was also taken.  The observations were completed as planned, and no anomalies were noted.
 
    1.2 Completed Two Sets of STIS/CCD 8837 (CCD Dark Monitor-Part 1)
 
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks.  The proposal completed with no reported problems.
 
    1.3 Completed STIS/CCD 8631 (Bright Quasar Close Lensing Search II)
 
        The Space Telescope Imaging Spectrograph (CCD) was used to expand the Cycle 8 second generation HST snapshot survey of bright quasars, optimized to find lenses with component image separations < 1".  The observations completed nominally.
 
    1.4 Completed Two Sets of STIS/CCD 8562 (Probing the Large Scale Structure: Cosmic Shear Observations)
 
        The Space Telescope Imaging Spectrograph (CCD) was used to probe the distortion of light bundles from distant galaxies, looking at the statistical properties of the intervening inhomogeneous {dark} matter distribution.  The proposal completed nominally.
 
    1.5 Completed STIS/CCD 8838 (Bias Monitor – Part 1)
 
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns.  The proposal completed nominally.
 
    1.6 Completed STIS/CCD 8870 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation IV)
 
        The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted, parallel observations as part of a POMS test proposal.  The observations completed with no anomalous activity.
 
    1.7 Completed STIS/CCD 8228 (The Black Hole Versus Bulge Mass Relationship in Spiral Galaxies)
 
        The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the galaxy NGC-2964 in order to study the black hole versus bulge mass relationship in spiral galaxies.  A wavelength calibration image was also taken.  The observations were executed as scheduled, and no anomalies were noted.
 
    1.8 Completed STIS/CCD 8700 (A STIS Spectroscopic Snapshot Survey of 3CR Radio Galaxies : The Nature of the Unresolved Nuclei)
 
        The Space Telescope Imaging Spectrograph (CCD) was used to perform STIS snapshot long-slit spectroscopy of these nuclei over 2900- 10300 AA to classify them {QSO, NLRG, BLRG, BL Lac, ldots} and establish their AGN characteristics down to the lowest nuclear luminosities, while still within powerful 3C radio galaxies.  The primary goals are : {1} to test unification schemes by identifying previously undetected signatures of low-level AGN/QSO activity in their spectra in the form of weak, broad emission lines, featureless continua, and UV excess, {2} to constrain the FR I/BL Lac connection by measuring the relative contribution of thermal {starburst}, non-thermal {optical synchrotron}, and line emission, and {3} to perform line and continuum diagnostics to determine the dominant ionization mechanisms in the circumnuclear regions {tenths of arcseconds} and the role of dust extinction in the NLRGs and BLRGs in the form of geometrically thick or thin tori and foreground dust.  The proposal completed as planned, with no reported anomalies.
 
    1.9 Completed STIS/CCD/MA1 8695 (The Nature and Distribution of O VI Absorbers in the Vicinity of Galaxies)
 
        The Space Telescope Imaging Spectrograph (CCD and MA1) was to study PG1259+593 with the E140M echelle mode to substantially improve the statistical base of observed O VI absorbers {including weak lines} for comparison with the cosmological models and to constrain the physical conditions and abundances in these systems and their relationships with galaxies.  With these observations we will {1} measure the number of O VI absorbers per unit redshift {dN/dz} with a limiting equivalent width of W_Lambda ~25 mAngstrom , {2} examine whether the O VI absorption arises in photoionized, collisionally ionized, or multiphase gas, {3} estimate the absorber metallicities, and {4} study the dependence of the absorber properties on the proximity of luminous galaxies.  The observations completed nominally.
 
    1.10 Completed STIS/MA2 8843 (Cycle 9 MAMA Dark Measurements)
 
        The Space Telescope Imaging Spectrograph (MA2) was used to perform the routine monitoring of the MAMA detector dark noise.  The proposal completed nominally.
 
********************************************
 
The following scheduled WF/PC-2 proposals were not executed since the instrument was placed in safe at approximately 276/114752Z.
 
Three Sets of WF/PC-2 8826 (Cycle 9 Supplemental Darks pt 1)
WF/PC-2 8816 (Cycle 9 UV Earthflats)
WF/PC-2 8656 (The Hydrogen-Burning Limit in the Globular Cluster NGC 6397) WF/PC-2 8822 (Decontaminations and Associated Observations Pt. 1/4) WF/PC-2 8544 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II)
 
2.0 FLIGHT OPERATIONS SUMMARY:
 
    2.1 Guide Star Acquisitions:
      Scheduled Acquisitions: 8
Successful: 8
 
  Scheduled Re-acquisitions: 7
Successful: 7
 
    2.2 FHST Updates:
Scheduled: 12
Successful: 12
 
    2.3 Operations Notes:
 
        HST operations are being conducted from the Science Institute Mission Operations Room (SIMOR).
 
        Using ROP SR-1A, the SSR EDAC error counter was cleared twice.
 
        Per ROP NS-3, the NSSC-1 status buffer was dumped three times.
 
        SI C&DH errors were reset on four occasions as directed by ROP NS-5.
 
        A TTR was written when a command re-transmit was required at 278/1935Z while a NSSC-1 load was being uplinked.
 
        The revised computer memory loads that included provisions for the recovery of  WF/PC-2 were received in the SIMOR at 278/2315Z, and eventually approved by all parties.  The loads were uplinked to the HST as planned.  WF/PC-2 recovery is scheduled to begin at 279/1800Z, with the WF/PC-2 decontamination to run from approximately 280/0700Z until approximately 280/2000Z.  The first WF/PC-2 shutter usage is to begin at approximately 280/2300Z.
 
3.0 SIGNIFICANT FORTHCOMING EVENTS:
 
        Continuation of normal science observations and calibrations, including those of WF/PC-2.

SpaceRef staff editor.